OpenStudy (anonymous):

The rms speed of CH4 at T1 K and SO2 at T2 K are the same the ratio T1/T2 is?

OpenStudy (anonymous):

@ujjwal @UnkleRhaukus

OpenStudy (ujjwal):

can you calculate the ratio of masses of CH4 and SO2?

OpenStudy (anonymous):

16/64 = 1/4

OpenStudy (ujjwal):

And you have \[V_{rms}=\sqrt\frac{3RT}{M}\]. so, you should be able to do it now. Put values. Obtain two equations. Divide them and that's it.

OpenStudy (anonymous):

yup i got it 1/4

OpenStudy (anonymous):

thxxx

OpenStudy (ujjwal):

wc buddy!